You have credit card debt of $ 30 comma 000 that has an APR​ (monthly compounding) of 17 %. Each month you pay the minimum monthly payment only. You are required to pay only the outstanding interest. You have received an offer in the mail for an otherwise identical credit card with an APR of 11 %. After considering all your​ alternatives, you decide to switch​ cards, roll over the outstanding balance on the old card into the new​ card, and borrow additional money as well. How much can you borrow today on the new card without changing the minimum monthly payment you will be required to​ pay?  

​Note: Make sure all calculations are held to at least five​ (5) decimal places.

The original loan payment is ​$ ?. ​(Round to the nearest​ cent.)

Amount of total borrowing on the new card today can be ​$ ?. ​(Round to the nearest​ dollar.)
Answer #1

1.
=Debt*rate/12
=30000*17%/12=425

2.
=Debt*rate/12*12/new rate-Debt
=30000*17%/12*12/11%-30000=16363.63636